Differentiate the following functions with respect to x (i) sinh x tanh x (ii) Z x √ x sin t t dt.


1
Expert's answer
2021-06-01T06:28:09-0400

(i) "\sinh x\tanh x"

differentiating with respect to x

"=\tanh x(\cosh x)+\sinh x(sec h^2x)"

"=\sinh x+\sinh x(sech^2 x)"


(ii) "Zx\sqrt{x}\sin t \space tdt"

treating Z and t as constant and differentiating with respect to x

"=\dfrac{3}{2}x^{\frac{3}{2}-1}Z\sin t\space tdt"

"=\dfrac{3}{2}x^{\frac{1}{2}}Z\sin t\space tdt"